OpenStudy (anonymous):

Solve: 8x - 2(x + 1) = 7x + 8 Answer -10 10 6 -6 Solve: 8x - 2(x + 1) = 7x + 8 Answer -10 10 6 -6 @Mathematics

OpenStudy (diyadiya):

x=-10

OpenStudy (diyadiya):

8x-2x-2=7x+8 6x-2=7x+8 x=-8-2 x=-10

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

x=-10

OpenStudy (mikezack123):

x=-10